OpenStudy (anonymous):

What is the solution to the equation? 2(x – 7) = 10x + 18

OpenStudy (anonymous):

First, expand LHS, 2x-14 2x-14=10x+18 2x-10x=18+14 -8x=32 x=-4 got it?

OpenStudy (anonymous):

thank you :)